OpenStudy (anonymous):

Let f(x)=x-2 and g(x)= x2 -7x-9 find f(g(-1)) a.-21 b.-3 c.3 d.21

OpenStudy (anonymous):

to find f(g(-1)) you take g(x) and substitute the whole problem for all instances of 'x' in f(x); the problem becomes f(g(x)) = (2x-7x-9)-2. From there you can simplify it to f(g(x)) = (-5x-9) -9, then further to (f(g(x)) = -5x - 18. After you've combined the two equations, substitute -1 in for x to find f(g(-1)). -5(-1) - 18 = 5 - 18 = -13

OpenStudy (anonymous):

er, I messed the math up... 1 sec

OpenStudy (anonymous):

I'm coming up with -6 now, are you sure you typed the question correctly?

OpenStudy (anonymous):

yea I typed it in correctly.